How could accounting methods, depreciate equipment, and other long-term assets be used to manipulate net income and asset values reported in the financial statements?
Answer #1
Depreciation is charging/spreading the cost of the asset ,to the Income statements spanning the useful life of the asset ,the concept of which is accepatable under all accounting principles ,over the entire globe.
But different companies choose different methods of depreciation such as :
1.Straight-line method--which charges the asset's cost less salvage, equally over the useful life of the asset.Here, the depreciation rate applied is 1/No.of useful years of the asset.
2.Declining balance/written down value method--wherein , the asset's cost less salvage is depreciated at the specified rate, on the diminishing /carrying value of the asset , every year,till the entire cost is recovered or the asset sold/retired.
3. Double declining balance method-- here the depreciation rate is twice the straight line rate ,applied on the diminishing balance of the asset,every year--but only till the asset reaches its salvage value--as the depreciable value at the beginning is the cost of the asset & not reduced by the salvage value.
Explaining with figures,
Let us suppose an asset or equipment with initial cost as   $ 500000, with a useful life of 8 yrs.Let us also assume $ 10000 as the salvage value.
Depreciation under the above different methods:
1.Under the straight-line method-- depreciable value of the asset=asset's cost less salvage ie. 500000-10000= 490000,
& the depreciation rate applied is 1/8=12.5%
equally over the useful life of the asset, ie. 8
So, the depreciation amount will be 490000*12.5% or 490000/8= $ 61250
& the asset value in the yr.-end balance sheet will be 500000-61250= 438750
2.Under the Declining balance/written down value method--depreciable value is the same $ 490000 as above & assuming
& the depreciation rate applied is 1/8=12.5%
Depreciation amount will be the same 490000*12.5%= 61250 as above, but
in the second year depreciation expense will be (500000-61250)*12.5%= 54844
Thus amount charged as depreciation expense will keep decreasing year after year ,
So, the net income wil be more in each of the year , resulting in more income taxes to be paid.
But asset value in the balance sheet at year- ends will be more , than under the above method from Yr. 2 onwards, because of lesser accumulated depreciation--
3. Under the Double declining balance method--depreciable value of the asset is its cost, ie $ 500000
& the depreciation rate applied is (1/8)*2=25%
So the depreciation amount will be the 500000*25%= 125000 in the first year &
(500000-125000)*25%= 93750 & so on till $ 10000 is reached.
No depreciation is charged after this point(ie. Salvage amount)
Net income is less because more depreciation is charged in the initial years , resulting in lesser income tax being paid.
The asset value in the balance sheet at end of yr. 1 will be the lowest at 500000-125000= 375000
Thus, adoption of different methods to depreciate equipment and other long-term assets can be used to the advantage of the management so as to manipulate net income figures and asset values reported in the financial statements , according to their wishes --ie. as they want them to be projected , lesser or more.